OpenStudy (anonymous):

Starting salaries for accounting majors have a population mean of $50,000 and a population standard deviation of $4,000. What is probability of an individual starts at: Over $55,000 Between $48,000 and $55,000 Under $49,000.

OpenStudy (amistre64):

what tools do you hav eto work with? tabels, calculators? etc ...

OpenStudy (anonymous):

Anything.

OpenStudy (amistre64):

then use a normalcdf function

OpenStudy (amistre64):

normalcdf(low, high, mean, sd) for infinity i just use a large value like -9999 or 9999

OpenStudy (anonymous):

a) 0.1056, correct?

OpenStudy (amistre64):

the wolf isnt behaving for me ... ill have to work the z scores 55-50 = 5, z = 5/4 P(z>5/4) = .10565 so yeah thats about right

OpenStudy (amistre64):

the normalcdf is just another way to express: P(a<z<b)

OpenStudy (anonymous):

thanks, about to give another problem that's similar.

OpenStudy (amistre64):

oh good, then the solution process will be similar :)